You just purchased a share of SPCC for ​$99. You expect to receive a dividend of ​$66 in one year. If you expect the price after the dividend is paid to be $114​,what total return will you have earned over the​ year? What was your dividend​ yield? Your capital gain​ rate?
Answer #1

Return = Dividend + ( Price after dividend - price before dividend)/ price before dividend

= 66 + ( 114 - 99 ) /99

= 66 + 15.15

= 81.15%

Dividend yield = Dividend / Price before dividend

= 66 / 99

= 66.67%

Capital gain rate =( Price after dividend - Price before dividend ) / Price before dividend

= ( 114 - 99 ) / 99

= 15.15%